# No-hair and almost-no-hair results for static axisymmetric black holes and ultracompact objects in astrophysical environments

### v1

@@ III.2 Deforming a black hole from the inside
This situation is carefully analyzed in Appendix B of Geroch and Hartle (1982). Any deformation from the inside of Schwarzschild spacetime needs to be encoded in the presence of inhomogeneities along the initially homogeneous rod. Linear inhomogeneities $\lambda(z)$ along the rod translates into a behaviour $U=f(z)\log r+o\left(\log r\right)$ as we approach the horizon, with $f(z)\in(0,1)$.
@@
It needs to be bounded to have a regular geometry. In fact, we can recall that the surface gravity of a stationary black hole is constant on the horizon due to the first law of black hole dynamics Bardeen et al. (1973); Racz and Wald (1996). Now, boundedness of the surface gravity requires that the density of the rod is homogeneous. Any kind of inhomogeneous behaviour leads to a divergent surface gravity. Furthermore, it also implies that the Kretschmann scalar diverges. Let us show this explicitly. Given the leading behavior for $U=f(z)\log r$, as $r\rightarrow 0$ for $z\in(-M,M)$, we have that the radial equation for $V$ becomes:
@@
Since the second term is of the form $\sim r\log r$, it gives a subdominant contribution as we approach $r\rightarrow 0$ and the leading order behaviour of $V$ is similar to that of $U$, namely $V=f^{2}(z)\log r+\text{subdominant}$. If we plug this ansatz in the surface gravity and the Kretschmann scalar we find:
@@
[…]
We see that if $f(z)$ is different from a constant function with a value either 00 or $1$ we find a singular geometry at $r\to 0$. A possible interpretation is the following. The constant density rod in the ALP represents an auxiliary source and when analyzing the black hole geometry we realize that it does not map to any kind of matter content (the Einstein tensor identically vanishes on the horizon). The inhomogeneous profiles do not correspond to such auxiliary source and do indeed represent a real matter content that arises in a non-vanishing Einstein tensor. Given that it is static (non-violating energy conditions) matter on top of an event horizon, it gives rise to a singular behaviour which is what we see in Eqs. (43) and (44).
@@ III.3 Deforming a black hole from the outside
Let us consider now more general black holes that are locally in vacuum but do not necessarily extend the vacuum region all the way up to infinity. Geroch and Hartle Geroch and Hartle (1982), showed that every black hole solution with an event horizon whose topology is a two-sphere is such that the function $U(r,z)$ takes the same values at both extremes of the horizon i.e. $U(0,M)=U(0,-M)$ and we have that $U(r,z)$ has the same distributional character as its Schwarzschild counterpart $U_{S}$ on the event horizon, namely that of a infinitely thin homogeneous rod located on the $z$ axis between $-M$ and $+M$. Equivalently, this means that every local black hole solution can be represented as
@@
with $\Delta U$ any solution to Laplace equation that is analytic and regular, even at the horizon. Such a solution represents a deformed black hole due to an external gravitational field, with the $V$ function being obtained from direct integration. These are all the local black holes in vacuum [^1] Although they were found previously Israel and Khan (1964); Mysak and Szekeres (1966); Israel (1973), an exhaustive and systematic analysis was presented by Geroch and Hartle Geroch and Hartle (1982), including all the black holes including those with toroidal topology.. However, as we discussed at the beginning of the section, finding a local solution for $U$ is not the whole story, since we still need to ensure that $V$ has a non-singular behaviour.
The horizon in these coordinates is located in the $z$-axis, between $-M$ and $M$. Hence, the $z$-axis is disconnected and made of two pieces, $(-\infty,-M)\cup(M,\infty)$. Thus, it is not straightforward that we can ensure that $V(0,z)$ vanishes at both pieces of the axis even if we fix that it vanishes at one of the two connected pieces. The condition $U(0,M)=U(0,-M)$ precisely ensures that $V(0,z)$ vanishes on both connected pieces of the axis if it vanishes on one of them. Indeed, given that we have $U=U_{S}+\Delta U$, we can also perform an splitting of the form $V=V_{S}+\Delta V$ and we know that both $U_{S}$ and $V_{S}$ obey the vacuum equations. $\Delta V$ should obey the following equation (see Eq. (15))
@@
Hence, we can ensure that $V$ vanishes on both pieces of the axis as long as $U(0,M)=U(0,-M)$ (understanding this expression as a limit). This ensures that we can make the function $V$ regular on the whole axis taking it as the integral in Eq. (20).
The final step to check is asymptotic flatness, which requires that the system is in equilibrium. If the ALP problem of the black hole surrounded by some matter content is in equilibrium, it leads to a solution that is regular and can be extended as a vacuum solution all the way up to infinity, with the matter located only on compact regions.
@@
Refer to captionRefer to caption
Figure 6: Here we depict the situations that we discuss in the main text. The shadowed black region represents the black hole, and the blue ring can be located either in the same plane as the black hole, or in a plane that is not the one of the black hole. In the former case we find that the integral vanishes, whereas in the latter case we find a finite result.
@@
Figure 7: Here we represent the integrand of (25) a function of $r$ for some specific choices of the parameters, namely we are depicting the value for $z=3$ and we are taking the radius and mass of the ring to be $M_{\text{ring}}=1,r_{\text{ring}}=1$ and the black hole mass $M=1$ The red curve corresponds to the integrand when the ring and the black hole are on the same plane, whereas the blue one corresponds to them lying on different planes, namely five units displaced in the z-axis. A priori, there is no apparent symmetry reason to expect a cancellation of the integral.
More in general, we can take any kind of matter content in equilibrium surrounding the black hole and consider a deformation which is given directly by solving the ALP with the suitable boundary conditions that correspond to the matter content decided. In fact, the boundary conditions associated with the matter content can be replaced with an effective density $\rho_{\text{eff}}(\bm{x})$ that we can integrate with the Green function $G(\bm{x},\bm{x^{\prime}})=1/4\pi\absolutevalue{\bm{x}-\bm{x^{\prime}}}$ | to find the contribution to the $U$ function given by the matter content we have:
$\Delta U(\bm{x})=\int d^{3}x^{\prime}\frac{\rho_{\text{eff}}(\bm{x^{\prime}})}{\absolutevalue{\bm{x}-\bm{x^{\prime}}}}.$ (49)
We find that as long as there is no density in the horizon, we are avoiding the coincidence limit in the Green function and no singular behaviour arises.
@@ IV Black holes in astrophysical environments
Consider a homogeneous rod source for the ALP located at $r=0$ between $z=M$ and $z=-M$ in the coordinate system that we are using. Furthermore, consider a set of compact toroidal two-surfaces $S^{i}$ enclosing the external regions containing matter, see Fig. 8 for a pictorial representation. Inside these surfaces the vacuum equations are not obeyed and a careful analysis of the matter content would be required to determine the metric inside. However, to analyze their influence outside, we can replace the matter content with a Dirichlet boundary condition [^3] This is similar to the analysis of systems like the electromagnetic field between mirrors, where a perfectly reflecting boundary condition is imposed as a way to model in a simple way the interaction between the field and the matter content in the boundary. In this way, we can focus only on the determination of the field from its equations of motion supplemented with this boundary condition..
@@
Figure 8: Schematic representation of the setup. The $z$-axis is depicted as a black solid line, with a dashed segment between $z=-M$ and $z=M$ where the event horizon is located. Finally, we have illustrated some compact surfaces enclosing matter content (they would correspond to topological torus once we consider the $\varphi$ direction too) and we replace the potential matter content inside with a boundary condition at the surface of the object.
@@
As we have shown in the previous section, the central black hole can only be represented by a homogeneous rod source with a single free parameter: its length or mass parameter $M$. Knowing this internal source, the value of the function $U$ at the toroidal two-surfaces, and the asymptotic condition that $U\to 0$ at infinity to ensure asymptotic flatness, we can solve the ALP equation for all vacuum regions in spacetime. In fact, we have a well-posed elliptic problem with Dirichlet boundary conditions, i.e., we know the harmonic function $U$ at the horizon, at the surfaces $S^{i}$ and we know that it goes to zero at infinity as a polynomial in $1/r$. Hence, the solution to the Laplace equation exists and is unique outside the surfaces by virtue of the classic theorems Gilbarg and Trudinger (2001).
@@
The final ingredient that we need to ask for is consistency, i.e., that the equilibrium condition that Eq. (25) is obeyed everywhere. This is required so that the system does not display any kind of singular (cosmic-string-like) behavior, and we are really representing a black hole configuration in a gravitational environment. The specific form of the function $\Delta U=U-U_{S}$ determines the specific departure from spherical symmetry of the distorted black hole.
Putting everything together, we have shown:
@@
    That given a gravitational environment, the horizon of a black hole departs from spherical symmetry in a unique and specific manner dictated by the environment. Any deviation from this natural shape will result in an irregularity of the geometry.
@@
    That any spheroidal shape for the horizon is in principle possible; one only needs to find which would be the necessary gravitational environment.
@@
    That deformations of the horizon generated by external matter content are mild in terms of curvature, in the sense that they give rise to bounded curvature contributions at the horizon.
@@
    That, in any case, for a solution with black holes and external matter to be regular everywhere, we need the equilibrium condition (25) in the ALP.
@@ V Almost-no-hair results for horizonless ultracompact objects
In the static axisymmetric case we have shown that we can find a uniqueness result under complete generality, i.e. without some of the conditions in Israel theorem (which appear necessary in the general static case). Let us now consider that instead of black holes we have horizonless ultracompact objects. By that we mean material objects whose first equipotential surface outside matter has a very small redshift value, $e^{2U_{\rm surf}}=\epsilon$, but not zero: $0<\epsilon\ll 1$.
One could think that if no horizon is present, any ultracompact object is possible, for instance compact but finite Curzon objects, thin rings, and non-spherically symmetric ultracompact spheroid objects. However, we can devise a new form of no-hair results by requiring that the Kretschmann scalar does not acquire trans-Planckian values, i.e. ${\cal K}<{\cal K}_{P}=1/\ell_{P}^{4}$ with $\ell_{P}$ representing the Planck length (or, more generally, assuming that all curvature invariants are bounded as proposed in the limiting curvature hypothesis Markov (1982, 1987); Frolov et al. (1990); Frolov and Zelnikov (2022); Frolov (2022)). This condition replaces the condition that the horizon is regular in the standard theorems. As we are going to show this condition puts strong constraints on the type of geometries that one can construct, in particular in the strength of the distortions from spherically symmetric stellar spacetimes that one can generate. The idea is that, on the one hand, we should not trust GR anyway when such large curvatures appear. On the other hand, one can expect that the dynamics of any system will try to avoid reaching configurations of huge curvatures if there are other possibilities with smaller curvatures available. Let us separately analyze the three situations we can have for static and axisymmetric exotic compact objects.
@@
When dealing with metrics close to the Curzon family we can think of a point-like region in Weyl coordinates but now with a finite extent. Outside a very small radius $R=R_{0}$ encompasing the object, the geometry can be defined by a specific combination of multipolar contributions. As described above, the Kretschmann scalar always grows unboundedly in some concrete directions depending on the specific multipolar expansion of the configuration.
Let us start with the simplest case, analyzing the monopole contribution. Given a fixed value of $\epsilon$ representing the minimum redshift attained by the configuration, we have a relation between $M^{(0)}$, $R_{0}$ and $\epsilon$ of the form
@@
The ratio $M^{(0)}/R_{0}$ must be very large in order to have $\epsilon\ll 1$. We can express $M^{(0)}$ in terms of $\epsilon$ and $R_{0}$ as before:
$M^{(0)}=\frac{1}{2}R_{0}\absolutevalue{\log\epsilon}.$ | . (51)
@@
where we have neglected subdominant terms which are smaller for small $R_{0}$. The limitation on the maximum value of the Kretschmann amounts to an inequality of the form
@@
$\frac{R_{0}}{\ell_{P}}\gtrsim\frac{1}{\sqrt{2}}\absolutevalue{\log\epsilon}^{3/2}\exp\left(\frac{1}{8}|\log\epsilon|^{2}\right),$ (54)
after substituting the specific function $\mathcal{R}_{0}\left(R,\pi/2;M^{0}\right)$ from Eq. (32). We are taking $\theta=\pi/2$ which is the direction in which the Kretschmann grows faster as we approach $R\rightarrow 0$. As a conclusion, a monopolar Curzon object with a minimum redshift surface (i.e. small $\epsilon$) must be very large (i.e. $R_{0}\gg\ell_{P}$) with an even larger $M^{(0)}$ if one does not want to have large curvatures. We must have the following separation of scales $M^{(0)}\gtrsim R_{0}\gg\ell_{P}$. Whereas the separation between $M^{(0)}$ is logarithmical in $\epsilon$ and hence not so large, see Eq. (51); the separation between $R_{0}$ and $\ell_{P}$ is much larger, see Eq. (54).
As we have already mentioned, an ultracompact monopolar Curzon object is actually not spherically symmetric (its apparent sphericity is merely an artifact of the Weyl coordinates). In fact it could be seen as representing the collapse of a cloud of matter tending towards a very big and thin lenticular object (see the illustrative picture in Fig. 9). It corresponds to an anisotropic collapse in which matter is splashed in one direction while it spreads over in the other two. In other words, this means that we need the extended object to be extremely flattened to avoid trans-Planckian curvatures.
@@
An equivalent discussion applies to any other ultracompact object based on a different multipolar combination. The difference now is that equipotential surfaces are no longer precise spheres in Weyl coordinates, resulting in objects that can have very complicated shapes.
@@
Given the convoluted expression for the Kretschmann scalar that we find for the ring, we will do an analysis based on purely dimensional grounds and inspecting the leading orders of the divergences in the potential and Kretschmann scalar. Qualitatively speaking the analysis is very similar to the previous case.
In the case of rings, the equipotential surfaces close to the ring are torii. For a given value of the mass $M$, the minimum value of the redshift $\epsilon$ is acquired for a torus of given radius $d$. In reality there is a dependence on the direction from which one approaches the center of the ring, but we eliminate this dependence here, considering $d$ an average distance to the center of the regular ring. We can understand this radius as the thickness of the ultracompact ring under consideration. The leading order of the redshift function imposes a relation between the mass of the ring and such radius:
$\epsilon\sim\exp(-M/d);\quad\longrightarrow\quad M\sim d\times\absolutevalue{\log\epsilon}.$ | . (55)
We can think that the mass of the ring is fixed by $\epsilon$ and $d$. Now, the trans-Planckian limitation is telling us that
${\cal K}\sim e^{N(M/d)^{n}}\times{\cal R}^{-4}<{\cal K}_{P}=\frac{1}{\ell_{P}^{4}},$ (56)
where we have substituted the Kretschmann by its leading order expansion expected on dimensional grounds. The length scale ${\cal R}$ has a very convoluted form. Inserting the previous expression for $M$ as a function of $d$ and $\epsilon$ we will obtain a condition of the form
$d>\ell_{P}F\times\absolutevalue{\log\epsilon},$ (57)
with $F$ a growing function with the argument |$\absolutevalue{\log\epsilon}$ | which tends to infinity as $\epsilon\to 0$. At the end, we find again a separation of scales $M\gtrsim d\gg\ell_{P}$.

### v3

@@ III.2 Deforming a black hole from the inside
This situation is carefully analyzed in Appendix B of Geroch and Hartle (1982). Any deformation from the inside of Schwarzschild spacetime needs to be encoded in the presence of inhomogeneities along the initially homogeneous rod, i.e. a function $\lambda(z)$. It translates into a behavior $U=f(z)\log r+o\left(\log r\right)$ as we approach the horizon $r\to 0$ with a fixed $z\in(-M,M)$. Here $f(z)=2\lambda(z)$. This is easy to understand, as when one approaches a segment closely enough, the local linear density dictates the leading behavior of the potential.
@@
It needs to be bounded to have a regular geometry. In fact, we can recall that the surface gravity of a stationary black hole is constant on the horizon due to the first law of black hole dynamics Bardeen et al. (1973); Racz and Wald (1996). Now, boundedness of the surface gravity requires that the density of the rod is homogeneous. Any kind of inhomogeneous behavior leads to a divergent surface gravity. Furthermore, it also implies that the Kretschmann scalar diverges. Let us show this explicitly. Given the leading behavior for $U=f(z)\log r$, as $r\rightarrow 0$ for $z\in(-M,M)$, we have that the radial equation for $V$ becomes:
@@
Since the second term is of the form $\sim r\log r$, it gives a subdominant contribution as we approach $r\rightarrow 0$ and the leading order behavior of $V$ is similar to that of $U$, namely $V=f^{2}(z)\log r+\text{subdominant}$. If we plug this ansatz in the surface gravity and the Kretschmann scalar we find:
@@
$+\text{finite terms}.$ (43)
We see that if $f(z)$ is different from a constant function with a value either 0 or $1$ we find a singular geometry at $r\to 0$. A possible interpretation is the following. The constant density rod in the ALP represents an auxiliary source and when analyzing the black hole geometry we realize that it does not map to any kind of matter content (the Einstein tensor identically vanishes on the horizon). The inhomogeneous profiles do not correspond to such auxiliary source and do indeed represent a real matter content that arises in a non-vanishing Einstein tensor. Given that it is static (non-violating energy conditions) matter on top of an event horizon, it gives rise to a singular behavior which is what we see in Eqs. (42) and (43).
@@ III.3 Deforming a black hole from the outside
Let us consider now more general black holes that are locally in vacuum but do not necessarily extend the vacuum region all the way up to infinity. Geroch and Hartle Geroch and Hartle (1982), showed that every black hole solution with an event horizon whose topology is a two-sphere is such that the function $U(r,z)$ takes the same values at both extremes of the horizon i.e. U(0,M)=U(0,M)U(0,M)=U(0,-M) and we have that $U(r,z)$ has the same distributional character as its Schwarzschild counterpart $U_{S}$ on the event horizon, namely that of a infinitely thin homogeneous rod located on the $z$ axis between M-M and +M+M. Equivalently, this means that every local black hole solution can be represented as
@@
with $\Delta U$ any solution to Laplace equation that is analytic and regular, even at the horizon. Such a solution represents a deformed black hole due to an external gravitational field, with the function $V$ being obtained by direct integration. These are all the local black holes in vacuum [^1] Although they were found previously Israel and Khan (1964); Mysak and Szekeres (1966); Israel (1973), an exhaustive and systematic analysis was presented by Geroch and Hartle Geroch and Hartle (1982), including all the black holes including those with toroidal topology.. However, as we discussed at the beginning of the section, finding a local solution for $U$ is not the whole story, since we still need to ensure that $V$ has a non-singular behavior.
The horizon in these coordinates is located in the $z$-axis, between M-M and $M$. Hence, the $z$-axis is disconnected and made of two pieces, $(-\infty,-M)\cup(M,\infty)$. Thus, it is not straightforward that we can ensure that $V(0,z)$ vanishes at both pieces of the axis even if we fix that it vanishes at one of the two connected pieces. The condition U(0,M)=U(0,M)U(0,M)=U(0,-M) precisely ensures that $V(0,z)$ vanishes on both connected pieces of the axis if it vanishes on one of them. Indeed, given that we have $U=U_{S}+\Delta U$, we can also perform an splitting of the form $V=V_{S}+\Delta V$ and we know that both $U_{S}$ and $V_{S}$ obey the vacuum equations. $\Delta V$ should obey the following equation (see Eq. (15))
@@
Hence, we can ensure that $V$ vanishes on both pieces of the axis as long as U(0,M)=U(0,M)U(0,M)=U(0,-M) (understanding this expression as a limit). This ensures that we can make the function $V$ regular on the whole axis taking it as the integral in Eq. (20).
The final step to check is asymptotic flatness, which requires that the system is in equilibrium. If the ALP of the black hole surrounded by some matter content is in equilibrium, it leads to a solution that is regular and can be extended as a vacuum solution all the way up to infinity, with the matter located only on compact regions.
@@
Figure 6: The shadowed black region represents the black hole, and the blue ring can be located either in the same plane as the black hole, or in a plane that is not the one of the black hole. In the former case we find that the integral vanishes, whereas in the latter case we find a finite result.
@@
Figure 7: Here we represent the integrand of (25) a function of $r$ for some specific choices of the parameters, namely we are depicting the value for $z=3$ and we are taking the radius and mass of the ring to be $M_{\text{ring}}=1,r_{\text{ring}}=1$ and the black hole mass $M=1$ . The red curve corresponds to the integrand when the ring and the black hole are on the same plane, whereas the blue one corresponds to them lying on different planes, namely five units displaced in the z-axis. A priori, there is no apparent symmetry reason to expect a cancellation of the integral.
More in general, we can take any kind of matter content in equilibrium surrounding the black hole and consider a deformation which is given directly by solving the ALP with the suitable boundary conditions that correspond to the matter content decided. In fact, the boundary conditions associated with the matter content can be replaced with an effective density $\rho_{\text{eff}}(\bm{x})$ that we can integrate with the Green function $|G(\bm{x},\bm{x^{\prime}})=1/4\pi\absolutevalue{\bm{x}-\bm{x^{\prime}}}$ to find the contribution to the $U$ function given by the matter content:
$|.\displaystyle\Delta U(\bm{x})=\int d^{3}x^{\prime}\frac{\rho_{\text{eff}}(\bm{x^{\prime}})}{\absolutevalue{\bm{x}-\bm{x^{\prime}}}}$. (48)
We find that as long as there is no density in the horizon, we are avoiding the coincidence limit in the Green function and no singular behavior arises.
@@ IV Black holes in astrophysical environments
Theorem: Given a regular (in the sense precised below) gravitational environment, there exists only one static, axisymmetric and asymptotically flat geometry containing a black hole which is non-singular. Moreover, the horizon of the black hole will depart from spherical symmetry (the shape when the environment is trivial) in a unique and specific manner that is completely dictated by the environment.
@@
Consider a homogeneous rod source for the ALP located at $r=0$ between $z=M$ and z=Mz=-M in the coordinate system that we are using. Furthermore, consider a set of compact and $\mathcal{C}^{\infty}$ toroidal two-surfaces [^3] While it may be possible to consider less regular surfaces, we assume them to be arbitrarily smooth for simplicity in our analysis. $S^{i}$ enclosing the external regions containing matter, see Fig. 8 for a pictorial representation. Inside these surfaces the vacuum equations are not obeyed and a careful analysis of the matter content would be required to determine the metric inside. However, to analyze their influence outside, we can replace the matter content by the value of the function $U$ induced on those surfaces. We restrict ourselves here to situations in which the functions $U$ at these surface are at least $\mathcal{C}^{2}$. [^4] Although it must be possible to relax this condition, this is enough for our purposes. Notice that we are not only demanding that the $U$ restricted to the surface is $\mathcal{C}^{2}$ but also across the surface, which automatically excludes the presence of any distributional matter content, i.e. thin shells. Then, the values of $U$ at the different toroidal surfaces act as Dirichlet boundary conditions. [^5] This is similar to the analysis of systems like the electromagnetic field between mirrors, where a perfectly reflecting boundary condition is imposed as a way to model in a simple way the interaction between the field and the matter content in the boundary. In this way, we can focus only on the determination of the field from its equations of motion supplemented with this boundary condition.
@@
Figure 8: Schematic representation of the setup. The $z$-axis is depicted as a black solid line, with a dashed segment between z=Mz=-M and $z=M$ where the event horizon is located. Finally, we have illustrated some compact surfaces enclosing matter content (they would correspond to topological torus once we consider the $\varphi$ direction too) and we replace the potential matter content inside with a boundary condition at the surface of the object.
@@
As we have shown in the previous section, the central black hole can only be represented by a homogeneous rod source with a single free parameter: its length or mass parameter $M$. Knowing this internal source, the value of the function $U$ at the toroidal two-surfaces, and the asymptotic condition that $U\to 0$ at infinity to ensure asymptotic flatness, we can solve the ALP equation for all vacuum regions in spacetime. In fact, we have a well-posed elliptic problem with Dirichlet boundary conditions, i.e., we know the harmonic function $U$ at the horizon, at the surfaces $S^{i}$, and we know that it goes to zero at infinity as a polynomial in $1/r$. Hence, the solution to the Laplace equation exists and is unique outside the surfaces by virtue of the classic theorems Gilbarg and Trudinger (2001).
@@
The final ingredient that we need to ask for is consistency, i.e., that the equilibrium condition (25) is obeyed everywhere. This is required so that the system does not display any kind of singular (cosmic-string-like) behavior, and we are really representing a black hole configuration in a gravitational environment. The specific form of the function $\Delta U=U-U_{S}$ determines the specific departure from spherical symmetry of the distorted black hole.
Putting everything together, we have shown that:
@@
    Given a gravitational environment, the horizon of a black hole departs from spherical symmetry in a unique and specific manner dictated by the environment. Any deviation from this natural shape will result in an irregularity of the geometry.
@@
    Any spheroidal shape for the horizon is in principle possible; one only needs to find which would be the necessary gravitational environment.
@@
    Deformations of the horizon generated by external matter content are mild in terms of curvature, in the sense that they give rise to bounded curvature contributions at the horizon.
@@
    In any case, for a solution with black holes and external matter to be regular everywhere, we need the equilibrium condition (25) in the ALP.
@@ V Almost-no-hair results for horizonless ultracompact objects
For static and axisymmetric black holes we have shown that we can find a uniqueness result under complete generality, i.e. without some of the conditions in Israel theorem, such as the constraint on the topology of the constant redshift surfaces and the finiteness of the area of the minimum redshift surface, which appear necessary in the general static case. Now, let us consider the scenario in which, instead of black holes, we are dealing with horizonless ultracompact objects. By this, we refer to material bodies whose first equipotential surface outside matter has an extremely high but finite redshift, characterized by a parameter $e^{2U_{\rm surf}}=\epsilon$, where $0<\epsilon\ll 1$.
Given that no-hair theorems require the existence of a horizon one could think that, if no horizon is present, any ultracompact object would be possible, for instance compact but finite Curzon objects, thin rings, and spheroidal but not spherically-symmetric ultracompact objects. However, a new form of no-hair results can be obtained by requiring that the Kretschmann scalar does not acquire trans-Planckian values Barceló et al. (2019), i.e. ${\cal K}<{\cal K}_{P}=1/\ell_{P}^{4}$ with $\ell_{P}$ representing the Planck length. More generally, this condition aligns with the limiting curvature hypothesis, which postulates that all curvature invariants should remain bounded Markov (1982, 1987); Frolov et al. (1990); Frolov and Zelnikov (2022); Frolov (2022). This curvature constraint effectively replaces the standard condition of a regular horizon in traditional no-hair theorems. As we will demonstrate, this constraint imposes stringent limitations on the permissible geometries, particularly in terms of the degree of distortion possible from spherical symmetry. The underlying physical motivation is that GR itself becomes unreliable when curvatures reach such extreme values. Indeed, this curvature constraint excludes not only non-spherical geometries but also spherically symmetric ultracompact configurations at the Planck scale from the set of plausible spacetime structures. This contrasts with the standard Israel theorem for black holes, which remains indifferent to the exact curvature values, as long as it remains bounded. In general, the GR dynamics is such that it tends to avoid configurations with extremely high curvatures except in cases where gravity becomes so strong that singularities inevitably form. Hence, it is reasonable to expect that any viable theory extending GR would naturally favor configurations with lower curvatures if such alternatives are accessible, thus avoiding configurations with extremely large curvatures.
As we saw before, in the static and axisymmetric case, there are only three possible sources in the ALP that can lead to arbitrarily high redshifts. Let us now analyze these three cases separately.
@@
When dealing with metrics close to the Curzon family we can think that these metrics describe the external region of an object with finite extent instead of a point-like region in Weyl coordinates. Outside a very small radius $R=R_{0}$ encompassing the object, the geometry can be defined by a specific combination of Weyl multipolar contributions. As described in Sec. III.1 above, the Kretschmann scalar always grows unboundedly in some concrete directions depending on the specific multipolar expansion of the configuration.
Let us start analyzing the simplest monopole contribution, namely $\ell=0$ in Eq. (36). Given a fixed value of $\epsilon$ representing the minimum redshift attained by the configuration, we have a relation between $M^{(0)}$, $R_{0}$, and $\epsilon$ of the form
@@
The ratio $M^{(0)}/R_{0}$ must be large in order to have $\epsilon\ll 1$. We can express $M^{(0)}$ in terms of $\epsilon$ and $R_{0}$ as:
$|.\displaystyle M^{(0)}=\frac{1}{2}R_{0}\absolutevalue{\log\epsilon}$. (50)
@@
where we have neglected subdominant terms for small $R_{0}$. The limitation on the maximum value of the Kretschmann amounts to an inequality of the form
@@
[…]
after substituting the specific function $\mathcal{R}_{0}\left(R,\pi/2;M^{0}\right)$ from Eq. (32). We are taking $\theta=\pi/2$ which is the direction in which the Kretschmann grows faster as we approach $R\rightarrow 0$. In conclusion, a monopolar Curzon-like object with a minimum redshift surface (characterized by a small $\epsilon$) must be extremely large (i.e., $R_{0}\gg\ell_{P}$), with an even larger $M^{(0)}$, to avoid high curvature values. This implies the following hierarchy of scales: $M^{(0)}>R_{0}\gg\ell_{P}$. While the separation between $M^{(0)}$ and $R_{0}$ depends logarithmically on $\epsilon$ and therefore does not need to be exceptionally large [see Eq. (50)], the separation between $R_{0}$ and $\ell_{P}$ must [see Eq. (53)].
As we have already mentioned, an ultracompact monopolar Curzon object is actually not spherically symmetric even if the redshift function $U$ is spherically symmetric. In fact, it could be seen as representing the collapse of a cloud of matter tending towards a very big and thin lenticular object (see the illustrative picture in Fig. 9). It corresponds to an anisotropic collapse in which matter is squeezed in one direction while it spreads over in the other two. However, our analysis shows that, given a mass for the lenticular object, it cannot be too flattened to avoid the presence of trans-Planckian curvatures at the lentil’s rim.
@@
An equivalent discussion applies to any other ultracompact object based on a different multipolar combination. The difference now is that equipotential surfaces are no longer precise spheres in Weyl coordinates, resulting in objects that can have more complex shapes.
@@
We have found a quite convoluted expression for the Kretschmann scalar using the software Mathematica with the help of the xAct package Martin-Garcia et al. . We have attached a Mathematica notebook containing the explicit expression which is not very illuminating. Thus, we have decided to perform an analysis based on purely dimensional grounds and inspecting the leading orders of the divergences in the potential and Kretschmann scalar. Qualitatively speaking the analysis is very similar to the previous case.
Close to the ring, the equipotential surfaces are tori. For a given value of the mass $M$, the minimum value of the redshift $\epsilon$ is acquired for a torus of given radius $d$. In reality, there is a dependence on the direction from which one approaches the center of the ring, but we ignore this dependence here, considering $d$ as an average distance to the center of the regular ring. We can understand this radius as the thickness of the ultracompact ring under consideration. The leading order of the redshift function imposes a relation between the mass of the ring and such radius:
$|.\displaystyle\epsilon\sim\exp(-M/d)\quad\Longrightarrow\quad M\sim d\absolutevalue{\log\epsilon}$. (54)
We can think that the mass of the ring is fixed by $\epsilon$ and $d$. Now, the bound on the curvature implies
${\cal K}\sim e^{N(M/d)^{n}}{\cal R}^{-4}<\frac{1}{\ell_{P}^{4}},$ (55)
[…]
with $F$ a growing function of its argument $|\absolutevalue{\log\epsilon}$ which tends to infinity as $\epsilon\to 0$. At the end, we would find again a separation of scales $M>d\gg\ell_{P}$.